OpenStudy (anonymous):

Solve the equation. 1/3x+9- 2/x+3=2

OpenStudy (anonymous):

answers: x=1 x=-23/6 x=-5/2 x=-5

OpenStudy (anonymous):

 Do the same thing that I did to get the common denominator.